### v1

# Diff-SSL-G-COMP: Towards a large-scale and diverse dataset for virtual analog modeling

Abstract
Virtual Analog (VA) modeling aims to simulate the behavior of hardware circuits via algorithms to replicate their tone digitally. Dynamic Range Compressor (DRC) is an audio processing module that controls the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production. In recent years, neural-network-based VA modeling has shown great potential in producing high-fidelity models. However, due to the lack of data quantity and diversity, their generalization ability in different parameter settings and input sounds is still limited. To tackle this problem, we present Diff-SSL-G-Comp, the first large-scale and diverse dataset for modeling the SSL 500 G-Bus Compressor. Specifically, we manually collected 175 unmastered songs from the Cambridge Multitrack Library. We recorded the compressed audio in 220 parameter combinations, resulting in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. Moreover, to facilitate the use of our proposed dataset, we conducted benchmark experiments in various open-sourced black-box and grey-box models, as well as white-box plugins. We also conducted ablation studies in different data subsets to illustrate the effectiveness of improved data diversity and quantity. The dataset and demos are on our project page: http://www.yichenggu.com/DiffSSLGComp/.
1 Introduction
Virtual Analog (VA) modeling aims to simulate analog audio devices digitally. Dynamic Range Compressor (DRC) is an audio processing module that compresses the dynamics of a track by reducing and amplifying the volumes of loud and quiet sounds, which is essential in music production [1]. VA modeling on DRC is important but always considered to be challenging due to its characteristics: non-linear, time-invariant, and long temporal dependency.
To model an analog compressor, early DSP-based methods utilized white-box models [1]. Although they can achieve outstanding performances, their model is usually device-specific and involves extensive human labor. As deep learning develops, neural-network-based black-box models [2, 3, 4, 5] have become popular in recent years due to their superior ability to model analog devices in a data-driven way, thus avoiding the high cost in human labor. More recently, following [6], grey-box models [7, 8] are also proposed, which combine the explainability and data-driven training from white and black-box models and achieve excellent performances.
Although these existing models have shown their potential in producing high-fidelity models, their generalization ability is still limited due to a lack of data quantity and diversity. In particular, existing VA modeling datasets [8, 9, 10, 11] primarily utilize synthetic test signals as inputs and only record scattered parameter combinations, which may significantly constrain the model performance when encountering real-world recordings and unseen parameters.
To tackle this issue, this work presents Diff-SSL-G-Comp, the first large-scale and diverse dataset for modeling the SSL 500 G-Bus Compressor  [^1] https://solidstatelogic.com/products/stereo-bus-compressor-module. Specifically, we manually selected 175 unmastered real-world songs from the Cambridge Multitrack Library  [^2] https://www.cambridge-mt.com/ms/mtk/ and recorded the compressed signals in 220 parameter combinations, which results in an extensive 2528-hour dataset with diverse genres, instruments, tempos, and keys. To facilitate the use of our dataset, we conducted benchmarking experiments on various open-sourced black-box and grey-box models, as well as available white-box plugins. We also conducted ablation studies on data subsets with different songs and data scales to illustrate the effectiveness of improved data diversity and quantity.
2 Related Work
Table 1: A comparison of Diff-SSL-G-Comp with existing VA modeling datasets regarding DRC.
Device Duration (hour) Type Parameters Range Combinations
UA 6176 Limiter [9] 0.66 Transistor-Based Limiter Attack Release Input Level Output Level Ratio 800 $\mu\text{s}$ 1100 ms 4 7 All 1
Ampeg Opto Comp [8] 3.61 Optical Compressor Compression Release Level [3, 10] [1, 10] s 6 5
Flamma FC21 [8] 3.61 Optical Compressor Comp EQ Volumn [1, 10] [1, 10] 10 5
Yuer RF-10 [8] 3.61 OTA Compressor Attack Sustain level [1, 10] ms [1, 10] ms 10 6
Teletronix LA-2A [10] 48.63 Optical Compressor Peak Reduction Switch Mode [0, 100] [Compressor, Limiter] 20
TubeTech CL-1B [11] 37.54 Optical Compressor Threshold Attack Release Ratio [-40, 0] dB [5, 300] ms [0.005, 10] s 1:[1, 10] 108
SSL 500 G-Bus-Comp (ours) 2528.53 VCA Compressor Threshold Attack Release Ratio [-40, 0] dB [0.1, 30] ms [0.1, 1.6] s 1:[1.5, 10] 220
DSP-based white-box models generally comprise a gain computer and a level detector with different detailed algorithm designs [1, 12], which has been well-studied over the years. Apart from this, works have also been proposed to increase computational efficiency; in particular,  [13] utilized block processing strategies for speeding up computation, and  [14] implemented the module via FFTs to use its parallel computing ability. Automatic DRC parameter control is another relevant research topic. Specifically,  [15] investigated automatic parameter control based on different extracted features, followed by  [16] to further extend to multi-track scenarios. With the development of machine learning techniques,  [17] proposed a regression model to control the parameters based on a referenced sound, followed by  [18] using a more potent random forest algorithm. Based on these white-box models, various studies have been proposed to simulate analog compressors, including creating virtual models for specific devices [19] or detailed components like optocouplers [20] and amplifiers [21]. In general, white-box models have high explainability in model design and good computational efficiency. However, due to the non-linear distortion in analog circuits, extensive human expert labor is usually needed to obtain a high-fidelity model in a specific device with long-time tuning. This makes it very expensive and impractical compared with the data-driven black-box and grey-box methods.
neural-network-based black-box models have been developed a lot in recent years.  [10] first proposed an autoencoder model to model various audio effects.  [2] utilized the LSTM model for optimizing the long-term dependencies, followed by [22] to further expand into the hyper RNN model with an in-detailed comparison between RNN, GRU, and LSTM models. To utilize the advantages of CNNs,  [23] first employed the WaveNet [24] structure on digital audio effects. Based on this work,  [3] proposed a temporal convolutional network (TCN) with larger receptive fields and huge dilation factors.  [4] further improved this architecture by integrating the feature-wise linear modulation (FiLM) [25] layers in modeling the parameter conditions. State Space Model (SSM) [26] is another technique to model long-term dependent time series via decomposing a dynamic system into structured state variables.  [5] first employed the S4 blocks in VA modeling, obtaining outstanding performances, followed by  [27] further adopting the latest S6 model [28]. With the development of differentiable digital signal processing (DDSP) [6], works are also proposed to utilize the DSP models’ explainability and efficiency. Specifically,  [29] proposed the differentiable biquad filters for machine learning applications, and  [30] integrated the biquad filter modeling with Koopman Networks [31] to operate in a higher dimensional state space. These advances have also made the neural grey-box models viable. In particular,  [7] utilized the classic white-box DRC [1] design with MLPs predicting the parameters in each time frame, followed by  [32] to further simplify the model into parametric Gains for compression and supplementary EQs for non-linear distortion.
Despite the rapid development of VA models, the publicly available datasets are still scarce, with limited data quantity and diversity. Table 1 illustrates the details of the existing datasets regarding DRC. Specifically, early attempts [9] primarily processed short instrument recordings in a specific parameter setting for nonparametric models. SignalTrain [10] presented the first parametric dataset in modeling the optical compressor LA-2A. It used various randomly generated test signals and a few instrument recordings as the input signals and recorded 20 equally sampled parameter combinations. After that,  [11] proposed the CL-1B dataset with real-world recordings as inputs and more parameter combinations. Recent works like [27] also presented datasets with more diverse devices but often with limited data scale and parameter combinations.
Data scaling has been shown effective in many areas [33, 34]. For instance, Mert [35] utilized a music mixture of 160K hours to scale up a self-supervised representation learning model with 330M parameters, obtaining outstanding performance in music information retrieval; Yue [36] constructed a 650K hours music mixture to train a 7B parameter model for music generation, obtaining state-of-the-art (SOTA) performance; Stable Audio [37] collected 73k hours of audio recordings, leading to SOTA audio generation model with 1B parameters; Emilia [38] presented a 101K hours open-sourced speech dataset, facilitating SOTA speech generations models [39, 40, 41]. Following these previous works, this paper presents Diff-SSL-G-Comp, the first large-scale and diverse dataset in VA modeling. We aim to address the limitations in existing datasets and explore the effectiveness of the improved data scale.
3 Diff-SSL-G-Comp
[…]
(a) Genre
[…]
(b) Instrument
Figure 1: Duration statistics (hours) of the unmastered songs used as input signals in Diff-SSL-G-Comp by genres and instruments.
As discussed in Section 2, existing VA modeling datasets are undiversified with limited data scales, which may restrict the model performance. To address this issue, we present Diff-SSL-G-Comp, an extensive and diverse dataset for modeling the SSL 500 G-BUS Compressor \@footnotemark. This section provides the construction details, statistics, and analysis of Diff-SSL-G-Comp.
3.1 Dataset Construction
Diff-SSL-G-Comp comprises unmastered songs with different genres, instruments, tempos, and keys processed with varying compression parameters. In particular, we manually selected 175 unmastered songs from the Cambridge Multitrack Library \@footnotemark. We used Reaper  [^3] https://www.reaper.fm/ as the Digital Audio Workstation (DAW) to process the data automatically. Specifically, we used the RME Fireface UFX+  [^4] https://rme-audio.de/fireface-ufx.html as the external audio interface and connected it to the ReaInsert. Then, we wrote a ReaScript to automatically send and receive signals from the hardware compressor via the audio interface. To match the level between the DAW and hardware compressor, we normalized all songs to -12 dB and applied a 5 dB input boost and a 5 dB output attenuation. We manually selected 144 widely used parameter combinations for processing after consulting six professional mastering engineers, which are: threshold [-28, -24, -20, -16], attack [0.1, 0.3, 1, 3], release [0.1, 0.4, 0.8, auto], ratio [2, 4, 10]. We additionally recorded 76 other randomly selected combinations as supplementary edge cases. All the audio data was recorded at a sampling rate of 44.1kHz.
Figure 2: The annotation pipeline of Diff-SSL-G-Comp. We utilized various pre-trained models to obtain information on each song’s key, tempo, genre, and instrument.
3.2 Dataset Statistics
We utilized various pre-trained models to annotate our data, as illustrated in Fig. 2. Specifically, we used the KeyCNN and TempoCNN model  [^5] https://github.com/hendriks73/directional_cnns proposed in [42] to obtain the global music tempo and key information. We split each song into a series of 10s segments and used the Qwen2-Audio [43]  [^6] https://huggingface.co/Qwen/Qwen2-Audio-7B to annotate each segment’s content, which will then be fed to a Llama3 [44]  [^7] https://huggingface.co/meta-llama model to organize and determine the genres and instruments of the whole song.
The statistical results of Diff-SSL-G-Comp on genres, instruments, tempos, and keys are illustrated in Fig. 1 and Fig. 3. From these results, we can conclude that:
  • The majority of genres in our dataset are Rock, Pop, Electronic, and Folk, with a small amount of other uncommon ones like Jazz and Hip Hop.
  • Most used instruments in our dataset are Bass, Drum, Guitar, Vocal, and Synth, with a considerable amount of Piano, String, Keyboard, and Brass. Niche instruments, like Woodwind, are also presented in the dataset.
  • Songs in our dataset are within the range of 70-160 beats per minute (BPM), and the majority of songs are distributed around 110-130 BPM.
  • Most songs in our dataset are in C, D, E, F, G, and A Majors, with a small number of remaining songs evenly distributed across other keys.
(a) Tempo
(b) Key
Figure 3: Tempo and Key statistics (occurrences) of the unmastered songs used as input signals in Diff-SSL-G-Comp.
(a) Acoustic Diversity
(b) Semantic Diversity
Figure 4: Comparison of acoustic and semantic diversities in input signals between Diff-SSL-G-Comp and the existing datasets. The plottings are obtained by applying the PCA algorithm to the SSL representations. We used MERT to extract acoustic embeddings and w2v-BERT 2.0 to extract semantic embeddings. For existing datasets, the compact cluster represents random noises, the diffused cluster represents test signals (sine, square, triangle waves, and their combinations), and the remaining scattered points represent real-world recordings.
3.3 Dataset Analysis
Unlike existing datasets, which primarily utilize noises and analysis signals, Diff-SSL-G-Comp comprises a collection of diversified real-world unmastered songs as the input signals. To quantify this diversity, we use self-supervised learning (SSL) models to investigate and compare their differences in acoustic and semantic feature spaces, following [38][45], and [46].
Specifically, to analyze the diversity of acoustic features, we leveraged a pre-trained MERT [35]  [^8] https://huggingface.co/m-a-p/MERT-v1-330M model to extract the acoustic representation (the 12th layer is used), which captures various acoustic characteristics such as timbre, style, key, etc. For the semantic diversity analysis, we employed a pre-trained w2v-BERT model [47]  [^9] https://huggingface.co/facebook/w2v-bert-2.0 to generate semantic representations (the last layer is used), capturing melody, lyrics, rhythm, etc. We then applied the Principal Component Analysis (PCA) algorithm to reduce the dimensionality of these representations to two. As illustrated in Fig. 4, most sample points in existing datasets are centered in two distant clusters, where the compact one represents the noise signals, and the diffused one represents the test signals (sine, square, triangle waves, and their combinations), and only a few points scattered aside, representing the real-world instrument recordings. Compared with the existing datasets, Diff-SSL-G-Comp exhibits a broader dispersion in the cluster representing real-world recordings, indicating richer acoustic and semantic characteristic coverage.
4 Experiments
Table 2: Benchmarking results of existing parametric black-box methods on seen and unseen test sets. The best and second best results of every column are bold and underlined.
System Configuration Condition #Params L1 ($\downarrow$) M-STFT ($\downarrow$)
Seen Unseen Seen Unseen
LSTM [2] 32 Channels Concat 5.0K 0.0290 0.0239 0.3954 0.4644
TVConcat 8.0K 0.0030 0.0028 0.3631 0.4523
96 Channels Concat 39.7K 0.0274 0.0237 0.4732 0.8123
TVConcat 45.7K 0.0028 0.0029 0.4256 0.5483
TCN [3] 5 Blocks 7 Kernel 4 Dilation FiLM 15.0K 0.0296 0.0251 0.5432 0.8647
TFiLM 42.0K 0.0066 0.0056 0.3755 0.4492
TTFiLM 17.3K 0.0271 0.0224 0.3903 0.4953
TVFiLM 17.7K 0.0252 0.0224 0.5957 0.9704
10 Blocks 3 Kernel 2 Dilation FiLM 20.1K 0.0088 0.0079 0.5158 0.6959
TFiLM 76.4K 0.0080 0.0067 0.3731 0.4427
TTFiLM 27.0K 0.0260 0.0215 0.3804 0.5057
TVFiLM 22.8K 0.0083 0.0069 0.3819 0.3983
GCN [4] 5 Blocks 7 Kernel 4 Dilation FiLM 29.0K 0.0271 0.0223 0.4760 0.5527
TFiLM 146.0K 0.0041 0.0034 0.3713 0.4045
TTFiLM 31.6K 0.0066 0.0024 0.3817 0.5766
TVFiLM 31.7K 0.0270 0.0226 0.3406 0.4147
10 Blocks 3 Kernel 2 Dilation FiLM 40.5K 0.0241 0.0200 0.6757 0.6346
TFiLM 278.0K 0.0267 0.0220 0.3497 0.4438
TTFiLM 48.0K 0.0063 0.0024 0.3549 0.5766
TVFiLM 43.2K 0.0272 0.0226 0.3238 0.4456
S4 [5] 4 Blocks 4 State Dimension FiLM 8.9K 0.0287 0.0246 0.8044 1.0532
TFiLM 30.0K 0.0277 0.0230 0.3576 0.4973
TTFiLM 10.2K 0.0030 0.0030 0.3884 0.4689
TVFiLM 11.6K 0.0283 0.0237 0.3898 0.5842
8 Blocks 32 State Dimension FiLM 29.7K 0.0103 0.0102 1.0552 1.2474
TFiLM 74.3K 0.0046 0.0043 0.4961 0.6098
TTFiLM 34.8K 0.0265 0.0225 0.4665 0.5898
TVFiLM 32.4K 0.0030 0.0031 0.3480 0.4930
Table 3: Benchmarking results of existing grey-box methods on seen and unseen test sets. The best and second best results of every column are bold and underlined.
System Signal Chain #Params L1 ($\downarrow$) M-STFT ($\downarrow$)
Static Gain Make-Up Gain Level Detector Seen Unseen Seen Unseen
GreyBoxDRC [7] Soft Knee Static Gain One-Pole 0.6K 0.0076 0.0066 1.0046 1.1312
Switching One-Pole 0.6K 0.0067 0.0072 0.8108 1.2388
RNN Mod. One-Pole 0.7K 0.0076 0.0070 1.0066 1.2251
Hard Knee GRU One-Pole 0.8K 0.0062 0.0074 1.1072 1.5134
Switching One-Pole 0.8K 0.0059 0.0061 0.8758 1.0888
RNN Mod. One-Pole 0.9K 0.0061 0.0070 1.1218 1.6492
ToneTwist [8] PEQ $\rightarrow$ Gain $\rightarrow$ PEQ $\rightarrow$ Gain 1.6K 0.0034 0.0034 0.4098 0.6004
PEQ $\rightarrow$ Phase Inversion $\rightarrow$ Gain $\rightarrow$ PEQ $\rightarrow$ Gain 2.0K 0.0200 0.0168 1.5964 1.4596
Table 4: Ablation results of the GCN model trained on different subsets on seen and unseen test sets. The best and second best results of every column in each setting are bold and underlined.
#Songs Duration (hour) L1 ($\downarrow$) M-STFT ($\downarrow$)
Seen Unseen Seen Unseen
100 0.05 0.0279 0.0262 0.4718 0.5699
0.5 0.0278 0.0226 0.3649 0.4838
5 0.0298 0.0222 0.3641 0.4539
50 0.0273 0.0227 0.3245 0.4520
500 0.0272 0.0226 0.3238 0.4456
5 50 0.0276 0.0231 0.4793 0.5876
10 0.0277 0.0230 0.4294 0.5159
25 0.0277 0.0227 0.3333 0.5030
50 0.0274 0.0224 0.3247 0.4502
100 0.0273 0.0227 0.3245 0.4520
In this section, we conducted benchmark experiments in various black-box and grey-box models and available commercial plugins to verify the effectiveness and facilitate the use of Diff-SSL-G-Comp. We also conducted ablation studies on different data subsets to illustrate the effectiveness of improved data scale and diversity.
4.1 Experiment Setup
4.1.1 Data Split and Processing
For the train and evaluation data split, we randomly selected 112 songs as the train set and used the remaining 63 songs as the test set. We used our manually selected 144 parameter combinations for training and the seen test distribution. The remaining 76 parameter combinations are used as the unseen test distribution to assess the models’ generalization ability in edge cases.
4.1.2 Training Schedules
All the models are trained using the AdamW [48] optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.999$, and a initial learning rate of 0.005. The ReduceLROnPlateau Scheduler is used with a factor of 0.5 and a patience of 10000 steps. All the experiments are conducted on a single NVIDIA H200 GPU with a batch size 16 and num workers of 16 for 500K steps. We use the Truncated Backpropagation Through Time (TBPTT) [49] with a 0.01s segment length (4410 samples) to reduce memory costs and enhance training efficiency while maintaining long-term dependencies.
4.1.3 Baselines and Configurations
We use the NablAFx [32] toolbox for conducting benchmarking experiments on baseline systems. Specifically, we use LSTM [2], TCN [3], GCN [4], and S4 [5] for black-box models. The LSTM model is conditioned on direct concatenation (Concat) or time-varying concatenation (TVConcat) [8]. The TCN, GCN, and S4 models are conditioned on FiLM [25], temporal FiLM (TFiLM) [50], tiny temporal FiLM (TTFiLM) [8], and time-varying temporal FiLM (TVFiLM) [8]. We use GreyBoxDRC [7] and two compressor simulation chains proposed in ToneTwist [8] for grey-box models with the original configurations. For commercial plugins, we utilize the available models from Solid State Logic [^10] https://store.solidstatelogic.com/plug-ins/ssl-native-bus-compressor-2, Softube [^11] https://www.softube.com/bus-processor, Overloud [^12] https://www.overloud.com/products/comp-g, and PSPaudioware [^13] https://www.pspaudioware.com/products/psp-busspressor.
4.1.4 Evaluation Metrics
We use the Amphion [51] toolkit for objective evaluation. We use the L1 and Multi-Resolution STFT loss to evaluate the time and frequency-domain errors following ToneTwist [8]. We additionally report the number of trainable parameters to show the model size.
4.2 Black-Box Methods
Table 2 illustrates the benchmarking results on black-box methods. It can be observed that 1) Regarding different model configurations, LSTM and TCN models will directly improve the model performance when increasing the model size, while only the models conditioned on TTFiLM and TVFiLM layers will improve in GCN and S4 models. We speculate this is because when these models are conditioned on the FiLM layer, the conditioning layer is not strong enough and will perform worse when increasing the model sizes. In contrast, the TFiLM layer is effective but has too many parameters, leading to learning problems when scaling up the model sizes. 2) Regarding different conditional layers, LSTM models with TVConcat perform significantly better than pseudo concatenation. In TCN, GCN, and S4 models, TVFiLM surprisingly performs the best, indicating the effectiveness of time-varying modeling when simulating analog compressors. It is also worth noting that the models conditioned with TFiLM generally perform second, followed by the TTFiLM models with similar performances, indicating its effectiveness in reducing model parameters and computational costs. 3) Regarding different model types, GCN generally performs the best, indicating the effectiveness of WaveNet-style dilated convolutions, followed by the S4 and TCN models conditioned on TTFiLM and TVFiLM. It is worth noting that the LSTM model with TVConcat outperforms many other baselines, showing the importance of the conditioning layer. 4) Regarding different test scenarios, LSTM models with TVConcat and TCN, GCN, and S4 models with TFiLM, TTFiLM, and TVFiLM achieve outstanding performance in both seen and unseen parameter settings. In contrast, a performance gap can be observed in simpler conditioning layers. This indicates the advanced generalization ability brought by the adequate temporal-varying conditional layers.
4.3 Grey-Box Methods
The benchmarking results on grey-box models are illustrated in Table 3. For pseudo-grey-box modeling, it can be observed that 1) Regarding different gain computer models, static gain with a soft knee generally performs better with different level detectors. This illustrates the model’s explainability since the analog module in the SSL-G-Bus compressor uses a soft knee where the knee width will automatically be computed by an internal algorithm based on the threshold and ratio  [^14] https://www.solidstatelogic.com/assets. 2) Regarding various level detectors, using the switching one-pole filter generally gives the best results, followed by using a simple one-pole filter. Using RNN-modulated one-pole filters, on the other hand, performs comparably worse. We speculate this is because, unlike the LA-2A compressor, where the optical circuit modules will bring a lot of non-linear distortion and coloration to the resulting audio, our VCA compressor uses simple level detectors; thus, switching to a more complex RNN-based model damages the performance. 3) Regarding different test sets, a comparable performance gap exists between the seen and unseen distributions. This is also because of the changing compressor curve in the analog module \@footnotemark, making it hard for grey-box models without feedback loops to learn that information. For analog simulation with effect chains, it is pretty surprising that using only two parametric gains for modeling the compression and two parametric EQs for modeling the non-linear distortions obtains the best performance — The fewer constraints give more improvement room for the non-linear deep learning modules. Experiments also show adding a simple phase inversion module would damage the model performance since there are no phasers in the actual analog module, confirming its effectiveness and explainability.
4.4 White-Box Plugins
To evaluate the development of academic NN-based models and further illustrate the effectiveness of our proposed dataset, benchmarking results on available white-box plugins are also reported, as shown in Table 4. It is worth noting that we do not conduct a comparative analysis between these plugins since they all come from different companies and are modeled in various environments, including using different analog devices from different manufacturers and using different audio interfaces, cables, and other hardware devices for optimization This will result in noticeable errors in evaluation, making the objective metric scores larger in multiple amounts. However, comparing these commercial plugins with existing academic black-box and grey-box models, a significant performance gap can still be observed, especially in extreme compression scenarios. Specifically, the plugin that comes from the PSPaudioware has an M-STFT score of 0.3047 and 0.2184 on the seen and unseen test sets, significantly outperforming the scores of 0.3238 and 0.3983 from the best academic NN-based models. This shows that the SOTA academic NN-based model in VA modeling still has a long way to go compared with industry plugins, which also illustrates the importance of our work since both model structure and datasets need to be improved for better performance.
4.5 Ablation Study
We also conducted ablation studies to illustrate the effectiveness of improving data quantity and diversity. We selected the GCN model conditioned with the TVFiLM layer as the baseline model and compared its performance when trained on different subsets. In particular, to control the data quantity, we fixed the number of total songs to 100 and control the length used to clip each song, resulting in 5 subsets from 3 minutes to 500 hours; to investigate the data diversity, we fixed the total data quantity to 50 hours and control the number of total songs with the adjusted clip lengths, resulting in 5 subsets from 5 songs to 100 songs. The detailed division and evaluation results are illustrated in Table 5. It can be observed that 1) increasing the data quantity steadily improves the model performance from 3 minutes to 500 hours, with the 50 hours as the division line for significant improvement, which is also confirmed by previous works [52]. 2) Increasing the data diversity is effective when there are only a few songs, and the improvement will be saturated until there are 50 different songs, especially in the unseen parameter settings. This is because different songs have different dynamics. Increasing the data diversity can help the model see more compression patterns, thus enhancing its generalization ability in unseen songs and parameter combinations.
5 Conclusion
In conclusion, this paper presents Diff-SSL-G-Comp, the first extensive and diverse dataset for DRC VA modeling. Our dataset comprises 2528 hours of processed unmastered songs in 220 parameter combinations with diverse genres, instruments, tempos, and keys. We provide benchmarking results on various open-sourced black-box and grey-box models, as well as available white-box plugins to facilitate the use of our dataset. We also provide ablation experiment results on different data subsets to illustrate the effectiveness of the improved data scale and quantity.
6 Acknowledgment
Table 5: Benchmarking results of existing commercial plugins on seen and unseen test sets. The best and second best results of every column are bold and underlined.
System L1 ($\downarrow$) M-STFT ($\downarrow$)
Seen Unseen Seen Unseen
Solid State Logic 0.0322 0.0175 0.4489 0.2943
Softube 0.0448 0.0237 0.7069 0.4546
Overloud 0.0326 0.0176 0.4738 0.3253
PSPaudioware 0.0269 0.0145 0.3047 0.2184
We acknowledge the computational resources provided by the Aalto Science-IT project. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call.